# Dark lens candidates from Gaia Data Release 3

### v1

@@
11institutetext: Astronomical Observatory, University of Warsaw, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 22institutetext: Las Cumbres Observatory, 6740 Cortona Drive, Suite 102, Goleta, CA 93117, USA 33institutetext: Department of Particle Physics and Astrophysics, Weizmann Institute of Science, Rehovot 76100, Israel 44institutetext: Research School of Astronomy and Astrophysics, Australian National University, Mount Stromlo Observatory, Cotter Road Weston Creek, ACT 2611, Australia 55institutetext: Zentrum für Astronomie der Universität Heidelberg, Astronomisches Rechen-Institut, Mönchhofstr. 12-14, 69120 Heidelberg, Germany 66institutetext: Institute of Theoretical Physics and Astronomy, Vilnius University, Saulėtekio al. 3, Vilnius, LT-10257, Lithuania 77institutetext: Center for Astrophysics and Cosmology, University of Nova Gorica, Vipavska 11c, SI-5270 Ajdovščina, Slovenia
Gravitational microlensing is a phenomenon that allows us to observe dark remnants of stellar evolution even if they no longer emit electromagnetic radiation. In particular, it can be useful to observe solitary neutron stars or stellar-mass black holes, providing a unique window through which to understand stellar evolution. Obtaining direct mass measurements with this technique requires precise observations of both the change in brightness and the position of the microlensed star and the European Space Agency’s Gaia satellite can provide both. We analysed events published in the Gaia Data Release 3 (Gaia DR3) microlensing catalogue using publicly available data from different surveys. Here we describe our selection of candidate dark lenses, where we suspect the lens is a white dwarf (WD), a neutron star (NS), a black hole (BH), or a mass-gap object, with a mass in a range between the heaviest NS and the least massive BH. We estimated the mass of the lenses using information obtained from the best-fitting microlensing models, the Galactic model and the expected distribution of the parameters. We found eight candidates for WDs or NS and two mass-gap objects.
@@
K. ${}^{,}$ Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3 kkruszynska@lco.global    Ł. Wyrzykowski Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    K. A. Rybicki Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    K. Howil Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. Jabłońska Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    Z. Kaczmarek Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    N. Ihanec Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. Maskoliūnas Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. Bronikowski Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    U. Pylypenko Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3
@@ 1 Introduction
There are still many outstanding questions connected with the remnants of stellar evolution. The most common stellar remnant is a white dwarf, and more than 95 per cent will become one by the end of their life (Fontaine et al., 2001). Our understanding of white dwarfs was expanded in recent years by Gaia and its superb parallaxes. The largest catalogue consists of over 350’000 high-confidence WD candidates, expanding almost ten times the amount of known WDs before Gaia (Gentile Fusillo et al., 2021). The best catalogue of known pulsars is two orders of magnitude smaller than the one we have for WDs in our Galaxy (Manchester et al., 2005). Our most limited observational material is on BHs, in particular solitary ones. Most of the known BHs are linked to binary systems found either through X-ray emission due to accretion of their companions (e.g. Corral-Santana et al., 2016) or as gravitational wave sources due to their merger (e.g. Abbott et al., 2019). Additionally, gravitational wave mergers are detected most frequently in distant galaxies. Recently, Shenar et al. (2022), El-Badry et al. (2023b), and El-Badry et al. (2023a) have reported on BH candidates also detected as non-interacting binary systems. However, the only known direct mass measurement for a solitary stellar-mass BH was recently presented for OGLE-2011-BLG-0462/MOA-2011-BLG-191 (Sahu et al., 2022; Lam et al., 2022; Mróz et al., 2022; Lam & Lu, 2023) using the gravitational microlensing phenomenon.
@@
where $\kappa=4G/c^{2}\mathrm{au}\approx 8.144\mathrm{mas}/M_{\odot}$, $\theta_{\mathrm{E}}$ is the angular Einstein Radius, which can be measured with astrometric microlensing, and $\pi_{\mathrm{E}}$ is the microlensing parallax obtained from modelling the time-series photometry. A combination of these two effects was used to detect a stellar-mass BH for the first time in Sahu et al. (2022) and Lam et al. (2022), who used astrometric observations from HST and photometric observations from the ground.
@@ 2.3 Data
The Wyrzykowski et al. (2023) catalogue was built using only Gaia $G$, $G_{\rm BP}$ and $G_{\rm RP}$ photometry, but for this work, we utilised data available from other surveys. In particular, we wanted to include information from microlensing surveys which have better cadence, especially in the Galactic bulge. We have cross-matched the Gaia sources with the OGLE survey (Udalski et al., 1992, 2015a). We have found 145 events in common with the public OGLE events. 130 events were published as a part of the OGLE-IV analysis of microlensing optical depth in the Galactic plane (Mróz et al., 2019, 2020). 78 events were published were also published as OGLE Early Warning System alerts (Udalski et al., 2015b), overlapping with the 130 events coming from the OGLE-IV papers. We downloaded all publicly available data. If the event was published in OGLE-EWS, Mróz et al. (2019) or Mróz et al. (2020), we used the data shared with the article. We performed a similar search with MOA survey (Abe et al., 1997; Bond et al., 2001) using its alert stream and we found 20 events in common. We found 32 events in common with KMTNet survey public alerts (Lee et al., 2014; Kim et al., 2016). Six events were published by Gaia Science Alerts. These events were published with preliminary photometry and without errors. We simulated the errors using the following formula, (Wyrzykowski et al., 2023):
@@
All data sources listed above were then used either at the preliminary or the detailed event modelling stages or both. We provide data used for this stage in a machine-readable online archive.
@@ 4 Detailed analysis of selected events
We conducted a case-by-case analysis of the 35 events selected in the previous step. We used MulensModel to generate the microlensing models and emcee (Foreman-Mackey et al., 2013) to explore the parameter space. In this step, we opted to use the KMTNet pySIS photometric data in magnitudes. For many events, we had to perform an outlier removal procedure. We have done this by fitting a preliminary PSPL model and removing all data points outside of the 3 $\sigma$ range of the residuals from the preliminary model. In some cases, we had to remove certain light curves, because they were too noisy. Table 3 provides the name of the fitted event and a list of data sets with the amount of data points for each light curve. Table 4 presents the median values of the posterior distributions (PDFs) obtained for best-fitting solutions. Here, we applied the following criteria to select the dark candidate events:
@@
This way we obtained 16 events, where at least one solution passed those criteria. For these events, we then estimated the lens distance and mass. From the remaining events, nine were better described by a non-parallax model, five didn’t pass the blending parameter criterion and for one, GaiaDR3-ULENS-057, we managed to only find a non-parallax solution.

### v2

@@
11institutetext: Astronomical Observatory, University of Warsaw, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 22institutetext: Las Cumbres Observatory, 6740 Cortona Drive, Suite 102, Goleta, CA 93117, USA 33institutetext: Department of Particle Physics and Astrophysics, Weizmann Institute of Science, Rehovot 76100, Israel 44institutetext: Research School of Astronomy and Astrophysics, Australian National University, Mount Stromlo Observatory, Cotter Road Weston Creek, ACT 2611, Australia 55institutetext: Zentrum für Astronomie der Universität Heidelberg, Astronomisches Rechen-Institut, Mönchhofstr. 12-14, 69120 Heidelberg, Germany 66institutetext: Institute of Theoretical Physics and Astronomy, Vilnius University, Saulėtekio al. 3, Vilnius, LT-10257, Lithuania 77institutetext: Center for Astrophysics and Cosmology, University of Nova Gorica, Vipavska 11c, SI-5270 Ajdovščina, Slovenia 88institutetext: Department of Physics, University of Warwick, Gibbet Hill Road, Coventry, CV4 7AL, UK 99institutetext: Institute for Space-Earth Environmental Research, Nagoya University, Nagoya 464-8601, Japan 1010institutetext: Department of Earth and Space Science, Graduate School of Science, Osaka University, Toyonaka, Osaka 560-0043, Japan $1$institutetext: Code 667, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA 1212institutetext: Department of Astronomy, University of Maryland, College Park, MD 20742, USA 1313institutetext: Institute of Natural and Mathematical Sciences, Massey University, Auckland 0745, New Zealand 1414institutetext: Department of Earth and Planetary Science, Graduate School of Science, The University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-0033, Japan 1515institutetext: Institute of Astronomy, Graduate School of Science, The University of Tokyo, 2-21-1 Osawa, Mitaka, Tokyo 181-0015, Japan 1616institutetext: Oak Ridge Associated Universities, Oak Ridge, TN 37830, USA 1717institutetext: Institute of Space and Astronautical Science, Japan Aerospace Exploration Agency, 3-1-1 Yoshinodai, Chuo, Sagamihara, Kanagawa 252-5210, Japan 1818institutetext: Sorbonne Universitè, CNRS, UMR 7095, Institut d’Astrophysique de Paris, 98 bis bd Arago, 75014 Paris, France 1919institutetext: Department of Physics, University of Auckland, Private Bag 92019, Auckland, New Zealand 2020institutetext: University of Canterbury Mt. John Observatory, P.O. Box 56, Lake Tekapo 8770, New Zealand
@@
K. Kruszyńska, Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3 kkruszynska@lco.global    Ł. Wyrzykowski Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    K. A. Rybicki Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    K. Howil Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. Jabłońska Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    Z. Kaczmarek Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    N. Ihanec Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. Maskoliūnas Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. Bronikowski Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    U. Pylypenko Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    A. Udalski Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    P. Mróz Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    R. Poleski Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    J. Skowron Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. K. Szymański Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    I. Soszyński Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    P. Pietrukowicz Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    S. Kozłowski Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    K. Ulaczyk Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    P. Iwanek Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. Wrona Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. Gromadzki Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    M. J. Mróz Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    F. Abe Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    K. Bando Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    R. Barry Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    D. P. Bennett Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    A. Bhattacharya Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    I. A. Bond Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    A. Fukui Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    R. Hamada Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    S. Hamada Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    N. Hamasaki Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    Y. Hirao Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    S. Ishitani Silva Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    Y. Itow Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    N. Koshimoto Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    Y. Matsubara Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    S. Miyazaki Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    Y. Muraki Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    T. Nagai Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    K. Nunota Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    G. Olmschenk Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    C. Ranc Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    N. J. Rattenbury Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    Y. Satoh Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    T. Sumi Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    D. Suzuki Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    P. J. Tristram Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    A. Vandorou Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3    H. Yama Dark lens candidates from Gaia Data Release 3Dark lens candidates from Gaia Data Release 3
@@
Gravitational microlensing is a phenomenon that allows us to observe dark remnants of stellar evolution even if they no longer emit electromagnetic radiation. In particular, it can be useful to observe solitary neutron stars or stellar-mass black holes, providing a unique window through which to understand stellar evolution. Obtaining direct mass measurements with this technique requires precise observations of both the change in brightness and the position of the microlensed star. The European Space Agency’s Gaia satellite can provide both. Using publicly available data from different surveys, we analysed events published in the Gaia Data Release 3 (Gaia DR3) microlensing catalogue. Here we describe our selection of candidate dark lenses, where we suspect the lens is a white dwarf (WD), a neutron star (NS), a black hole (BH), or a mass-gap object, with a mass in a range between the heaviest NS and the least massive BH. We estimated the mass of the lenses using information obtained from the best-fitting microlensing models, the source star, the Galactic model and the expected distribution of the parameters. We found eleven candidates for dark remnants: one WDs, three NS, three mass-gap objects, and four BHs.
@@ 1 Introduction
There are still many outstanding questions connected with the remnants of stellar evolution. The most common stellar remnant is a white dwarf, and more than 95 per cent will become one by the end of their life (Fontaine et al., 2001). Our understanding of white dwarfs was expanded in recent years by Gaia and its superb parallaxes. The largest catalogue consists of over 350’000 high-confidence WD candidates, expanding almost ten times the amount of known WDs before Gaia (Gentile Fusillo et al., 2021). The best catalogue of known pulsars is two orders of magnitude smaller than the one we have for WDs in our Galaxy (Manchester et al., 2005). Our most limited observational material is on BHs, in particular solitary ones. Most of the known BHs are linked to binary systems found either through X-ray emission due to accretion of their companions (e.g. Corral-Santana et al., 2016) or as gravitational wave sources due to their merger (e.g. Abbott et al., 2019). Additionally, gravitational wave mergers are detected most frequently in distant galaxies. Recently, Shenar et al. (2022), El-Badry et al. (2023b), and El-Badry et al. (2023a); Chakrabarti et al. (2023) have reported on BH candidates also detected as non-interacting binary systems. However, the only known direct mass measurement for a solitary stellar-mass BH was recently presented for OGLE-2011-BLG-0462/MOA-2011-BLG-191 (Lam et al., 2022; Sahu et al., 2022; Mróz et al., 2022; Lam & Lu, 2023) using the gravitational microlensing phenomenon.
@@
where $\kappa=4G/c^{2}\mathrm{au}\approx 8.144\mathrm{mas}/M_{\odot}$, $\theta_{\mathrm{E}}$ is the angular Einstein Radius, which can be measured with astrometric microlensing, and $\pi_{\mathrm{E}}$ is the microlensing parallax obtained from modelling the time-series photometry. A combination of these two effects was used to detect a stellar-mass BH for the first time in Lam et al. (2022) and Sahu et al. (2022), who used astrometric observations from HST and photometric observations from the ground.
@@ 2.3 Data
The Wyrzykowski et al. (2023) catalogue was built using only Gaia $G$, $G_{\rm BP}$ and $G_{\rm RP}$ photometry, but for this work, we utilised data available from other surveys. In particular, we wanted to include information from microlensing surveys which have better cadence, especially in the Galactic bulge. We have cross-matched the Gaia sources with the OGLE survey (Udalski et al., 1992, 2015a). We have found 145 events in common with the public OGLE events. 130 events were published as a part of the OGLE-IV analysis of microlensing optical depth in the Galactic plane (Mróz et al., 2019, 2020). 78 events were published were also published as OGLE Early Warning System alerts (Udalski et al., 2015b) [^2] https://ogle.astrouw.edu.pl/ogle4/ews/ews.html, overlapping with the 130 events coming from the OGLE-IV papers. We downloaded all publicly available data. If the event was published in OGLE-EWS, Mróz et al. (2019) or Mróz et al. (2020), we used the data shared with the article. We performed a similar search with MOA survey (Abe et al., 1997; Bond et al., 2001) using its alert stream and we found 20 events in common. We found 32 events in common with KMTNet survey public alerts (Lee et al., 2014; Kim et al., 2016) [^3] https://kmtnet.kasi.re.kr/~ulens/. Six events were published by Gaia Science Alerts [^4] http://gsaweb.ast.cam.ac.uk/alerts/home. These events were published with preliminary photometry and without errors. We simulated the errors using the following formula, (Wyrzykowski et al., 2023):
@@
All data sources listed above were then used either at the preliminary or the detailed event modelling stages or both. We provide data used for this stage in a machine-readable online archive [^5] https://github.com/KKruszynska/gdr3_dark_lenses.
@@ 4 Detailed analysis of selected events
We conducted a case-by-case analysis of the 35 events selected in the previous step. We used MulensModel to generate the microlensing models and emcee (Foreman-Mackey et al., 2013) to explore the parameter space. In this step, we used the KMTNet pySIS photometric data in magnitudes. For KMTNet data, when possible, we pre-processed data, removing any points that had a negative value of FWHM column or with photometric uncertainty larger than 1 magnitude. For OGLE and MOA events, we used re-processed data coming from the end-of-the-season DIA photometric reduction pipelines (Udalski et al., 2015a; Bond et al., 2001). We also applied a correction procedure for uncertainties following Skowron et al. (2016) to OGLE data from bulge fields. Finally, we re-scaled the photometric uncertainties using this formula:
$\sigma^{\prime}_{n,i}=k_{n}\sigma_{n,i},$ (3)
where $\sigma^{\prime}_{n,i}$ is the re-scaled $i$-th uncertainty of the $n$-th telescope’s light-curve, $k_{n}$ is the scale factor for the $n$-th telescope, and $\sigma_{n,i}$ is the original $i$-th uncertainty of the $n$-th telescope’s light-curve. We obtained the scale factors in the following manner:
  1. 1.
    we fitted preliminary PSPL models with and without parallax;
  2. 2.
    we selected a model with the smallest $\chi^{2}$ value;
  3. 3.
    using this model we used it as a starting point, we ran an MCMC fit with scale factors as one of the fit parameters.
If the scale factor of the median solution found in the final step exceeded 1.0 for a given telescope, we used this value and then we re-fitted the PSPL models with and without parallax. For events GaiaDR3-ULENS-003, GaiaDR3-ULENS-032, GaiaDR3-ULENS-118, GaiaDR3-ULENS-196, and GaiaDR3-ULENS-284 we had to use the PSPL without parallax model instead of the best model to find the scale factors. For event GaiaDR3-ULENS-025, GaiaDR3-ULENS-143, we couldn’t find scale factors due to poor event coverage. We report the values for each scale factor in LABEL:tab:fit_data. For some events, we had to perform an outlier removal procedure. We have done this before applying the uncertainty re-scaling. We used the best-fitting preliminary model (step 2 of the uncertainty re-scaling procedure). Then we removed all data points outside the 3 $\sigma$ range of the residuals from the preliminary model for a given light curve. We marked those light curves in bold in Table LABEL:tab:fit_data. In some cases, we had to remove certain light curves, because they were too noisy, or carried little information about the event. We marked those data sets with a strike-through text in Table LABEL:tab:fit_data. Table LABEL:tab:fit_data provides the name of the fitted event and a list of data sets with the amount of data points for each light curve. Table LABEL:tab:fit_best_model presents the median values of the posterior distributions (PDFs) obtained for best-fitting solutions. KMTNet data for event GaiaDR3-ULENS-047/KMT-2015-BLG-0157 revealed features around the peak, for which a PSPL model was insufficient in characterising, so it was excluded from further analysis. GaiaDR3-ULENS-284 had a large parallax value, which means other effects should be included. We excluded that event from further analysis. GaiaDR3-ULENS-057 had only three points at magnification, so we could only find a non-parallax solution. We named the solutions with the following convention: ”GaiaDR3-AAA-BC”, where AAA is the number assigned to the event in Wyrzykowski et al. (2023), B is a string of letters denoting which type of Gaia data was used: ”G” for events where we used the Gaia DR3 photometry, and ”GSA” for events where we opted for a Gaia Science Alerts lightcurve, finally C is a sign of the $u_{\mathrm{0}}$ of the solution (”+” for positive and ”-” for negative). If there was more than one solution with the same $u_{\mathrm{0}}$ sign, we numbered them starting from one.
To select the dark lens candidate sample, we applied the following criteria to the modelled events:
@@
This way we obtained 14 events, where at least one solution passed those criteria. For these events, we then estimated the lens distance and mass. From the remaining events, a non-parallax model better described two, seven didn’t pass the blending parameter criterion, and six did not pass the $\pi_{\rm E}$ criterion. We found four events that passed the blending and $\chi^{2}$ criterion but failed the $\pi_{\rm E}$ criterion in the calculated three sigma, however, their $\pi_{\rm E}$ distribution was inconsistent with 0. We analysed these solutions further but displayed their results in a different table.
5 Source stars
To determine the properties of the lens, we have to determine the properties of the source. Ideally, we would obtain the distance to the source, but that isn’t always possible. Instead, we decided to find the angular stellar radius of the source $\theta_{*}$ and use it as a prior during lens mass and distance estimation. We followed different procedures, depending on the event location and available information.
We could use the colour-magnitude diagrams (CMDs) calibrated to the OGLE-III data for events with MOA data (GaiaDR3-ULENS-035, GaiaDR3-ULENS-069, GaiaDR3-ULENS-073, GaiaDR3-ULENS-088, GaiaDR3-ULENS-155, GaiaDR3-ULENS-343, GaiaDR3-ULENS-353). First, we determined the red clump centre (RCG) location following the procedure outlined in Nataf et al. (2013). We use the de-reddened RCG distance modulus determined in Nataf et al. (2013) for each event [^6] https://ogle.astrouw.edu.pl/cgi-ogle/getext.py, and find the reddened distance modulus of the RCG using fitted position on the CMD and absolute magnitude of the RCG $M_{I,\mathrm{RCG}}=(-0.12,1.06)\leavevmode\nobreak\ \mathrm{mag}$ Nataf et al. (2013); Bensby et al. (2013). Then we calculate the extinction in $A_{I}$ and $A_{V}$ and use it to find the de-reddened magnitude and colour of the source. In the case of GaiaDR3-ULENS-353, the blending parameter was negative, so instead of using calculated source magnitude, we used the baseline magnitudes to determine source brightness and colour. Finally, we use these values to determine the angular stellar radius of the source star using relations from Adams et al. (2018).
Other sources were more difficult. If the source was located towards the Galatic Bulge, we assumed that the Bulge is located 8.1 kpc and extends for 2.4 kpc. We used this value to determine the de-reddened distance modulus to the red clump centre. We constructed a CMD in $V$ and $I$ data using Gaia DR3 sources. We selected sources within 30’ of the event, that had the Renormalized Unit Weight Error (RUWE) parameter smaller than 1.4, astrometric parallax error not larger than 20% of the measured value, and that had available GSP-Phot solutions (Andrae et al., 2023). Then, we transformed their colours into $V$ and $I$ bands following relations from Busso et al. (2022) Section 5.5.1777https://gea.esac.esa.int/archive/documentation/GDR3/Data_processing/chap_cu5pho/cu5pho_sec_photSystem/cu5pho_ssec_photRelations.html. We determined the position of the RGC following the procedure in Nataf et al. (2013), and used the calculated distance modulus to find the extinction in $I$ and $V$ bands. Finally, we found the $\theta_{*}$ using relation from Adams et al. (2018). We used this method for events GaiaDR3-ULENS-025, GaiaDR3-ULENS-089, GaiaDR3-ULENS-142, and GaiaDR3-ULENS-270. For many of these events, some information was missing. When baseline magnitude and blending in $G_{\rm BP}$ and/or $G_{\rm RP}$ filter, we used blending parameter in $G$ (for $G_{\rm BP}$) or in $I$ (for $G_{\rm RP}$) bands and Gaia DR3 entry for this source to determine the missing brightness for the $\theta_{*}$ estimation. We couldn’t use this method for events GaiaDR3-ULENS-331 and GaiaDR3-ULENS-363, because they were too dim compared to the data with GSP-Phot entries to infer the extinction.
Finally, for events located towards the Galactic disc, we couldn’t determine the de-reddened distance modulus towards the RCG, and therefore $\theta_{*}$. This affected events GaiaDR3-ULENS-103, GaiaDR3-ULENS-118, and GaiaDR3-ULENS-259.
All determined values can be found in Table LABEL:tab:source_prop.